OpenStudy (anonymous):

Liz works two part time jobs for a total of 34 hours a week. Job A pays $6.00 per hour and Job B pays her $7.20 per hour. How many hours did she work at each job the week that she made $226.80?

OpenStudy (anonymous):

a=amount of hours at job A b= amount of hours at job B then: a+b=34 (hours) 6a+7.2b=226.8 solve it for a &b - let me know what are your a& b

OpenStudy (anonymous):

I posted this before... what is the problem? can't solve the system? questions?

OpenStudy (anonymous):

lets say she works A hours at job A and B hours at job B together she works for a total of 34 hours. so A+B = 34 ------------------------- equation 1 she gets paid $6 at job A per hour. so, for A hours she gets paid 6A dollars similarly she gets paid 7.2B dollars for B hours in job B she earned 226.8 dollars totally, so 6A+7.2B = 226.8 ----------------- equation 2 now you have two equations and two unknowns, solve equations 1 and 2 for A and B

OpenStudy (anonymous):

lets try this. one job pays $6 per hour, and lets say she works x hours. so she makes 6x dollars. the other pays $7.20 per hour and if she works a total of 34 hours, then if she works x hours on job A she works (34-x) on job B. then her total amount is \[6x+7.20(34-x)\] and we know that the total is 226.80 so we set \[6x+7.20(34-x)=226.80\] and solve for x

OpenStudy (anonymous):

i would multiply by ten to clear the stupid decimals and get \[60x+72(34-x)=2268\] and then it is easier. \[60x+2448-72x=2268\] \[-12x=2268-2448=-180\] \[x=\frac{180}{12}=15\]